# Matrix-Transformation Based Low-Rank Adaptation (MTLoRA): A Brain-Inspired Method for Parameter-Efficient Fine-Tuning

### 12 Mar 2024 v1

@@ Abstract
Fine-tuning techniques based on Large Pretrained Language Models (LPLMs) have been proven to significantly enhance model performance on a variety of downstream tasks and effectively control the output behaviors of LPLMs. Recent studies have proposed numerous methods for fine-tuning a small number of parameters based on open-source LPLMs, reducing the demand for computational and storage resources. Among these, reparameterization fine-tuning methods represented by LoRA (Low-Rank Adaptation) have gained popularity. We find that although these methods perform well in many aspects, there is still considerable room for improvement in terms of complex task adaptability, performance, stability, and algorithm complexity. In response to this, inspired by the idea that the functions of the brain are shaped by its geometric structure, this paper integrates this idea into LoRA technology and proposes a new matrix transformation-based reparameterization method for efficient fine-tuning, named Matrix-Transformation based Low-Rank Adaptation (MTLoRA). The spatiotemporal patterns of brain neural activity are the excitation of different wavelength characteristic patterns of its geometric structure. MTLoRA aims to dynamically alter its spatial geometric structure by applying a transformation-matrix $T$ to perform linear transformations, such as rotation, scaling, and translation, on the task-specific parameter matrix, generating new matrix feature patterns (eigenvectors) to mimic the fundamental influence of complex geometric structure feature patterns in the brain on functions, thereby enhancing the model’s performance in downstream tasks. The transformation-matrix $T$ contains four different structures, each designed to simulate the geometric feature patterns of the brain at different levels. In Natural Language Understanding (NLU) tasks, it is evaluated using the GLUE benchmark test, and the results reveal that MTLoRA achieves an overall performance increase of about 1.0% across eight tasks and reduces the standard deviation by 0.7% in the Corpus of Linguistic Acceptability (CoLA) task; in Natural Language Generation (NLG) tasks, MTLoRA improves performance by an average of 0.95% and 0.31% in the DART and WebNLG tasks, respectively.
@@ 1 Introduction
In this paper, through empirical experiments on two categories of 11 tasks in NLU and NLG, the effectiveness of the MTLoRA method is verified. Compared to the LoRA method, it is evident from the experimental results table 1 that MTLoRA improves performance by about 1.54% ($\sigma$=0.1%) on the CoLA (Warstadt et al., 2019) task, effectively reducing the standard deviation; it enhances performance by about 3.61% ($\sigma$=0.8%) on the RTE (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) task; it increases performance by about 1.71% ($\sigma$=0.4%) on the MRPC (Dolan and Brockett, 2005) task; and it boosts performance by about 0.88% ($\sigma$=0.0%) on the QQP (Quora Question Pairs) task. From the experimental results table 2, it is known that MTLoRA achieves an average performance increase of about 0.95% ($\sigma$=0.1%) on the DART (Nan et al., 2020) task and about 0.31% ($\sigma$=0.3%) on the WebNLG (Gardent et al., 2017) task.
@@ 4.2.3 Experimental Results
$\mathbf{87.49}_{\pm.1}$ $\mathbf{94.5}_{\pm.1}$ $\mathbf{63.11}_{\pm.1}$ $\mathbf{91.25}_{\pm.0}$ $\mathbf{92.6}_{\pm.2}$ $\mathbf{87.73}_{\pm.8}$ $\mathbf{90.2}_{\pm 1.0}$ $\mathbf{91.37}_{\pm.0}$ $\mathbf{87.28}$
@@
In contrast, the ICFM structure shows robust performance across numerous tasks, particularly excelling in tasks related to semantic similarity and entailment when there is a lack of comprehensive corpus data. For example, it achieves an average performance boost of approximately 1.71% ($\sigma$=0.4%) on the MRPC task and 1.08% ($\sigma$=1.0%) on the RTE task.
@@ 4.3.3 Experimental results
MTLoRA(best) $70.1$ $48.31$ $55.98$
MTLoRA $69.58_{\pm.3}$ $\mathbf{48.05}_{\pm.1}$ $\mathbf{55.86}_{\pm.1}$
@@
MTLoRA-ICFM(best) $70.01$ $48$ $56$
MTLoRA-ICFM $69.57_{\pm.4}$ $47.87_{\pm.0}$ $55.82_{\pm.1}$
@@
The incorporation of the SHIM structure led to an enhancement in the model’s efficacy on the DART and WebNLG tasks, with an improvement of approximately 0.83% ($\sigma$=0.0%) and 0.56% ($\sigma$=0.1%), respectively. Similarly, the implementation of the ICFM structure resulted in a performance increment of roughly 0.77% ($\sigma$=0.0%) for DART and 0.54% ($\sigma$=0.1%) for WebNLG. The application of the CTCM structure contributed to a performance uplift on the DART and WebNLG tasks by approximately 0.88% ($\sigma$=0.0%) and 0.27% ($\sigma$=0.2%), respectively. Lastly, the adoption of the DTSM structure facilitated a performance boost on the DART and WebNLG tasks by an estimated 0.95% ($\sigma$=0.1%) and 0.31% ($\sigma$=0.3%), respectively.
@@ 4.4 Experimental Analysis
To verify whether the MTLoRA method actually reduces the model’s performance on the E2E task in NLG, we re-executed the performance test of LoRA on the E2E task using the same computational environment as MTLoRA. The experimental results show that LoRA achieves an average BLEU score of 69.52% on the E2E task, with a standard deviation of 0.5%. Its average BLEU score is comparable to that of MTLoRA, but with an increased standard deviation. This indicates that the experimental computational environment also has an impact on the results, suggesting that MTLoRA does not actually reduce the model’s performance on the E2E task in NLG.

### 18 Mar 2024 v2

@@ Abstract
Fine-tuning techniques based on Large Pretrained Language Models (LPLMs) have been proven to significantly enhance model performance on a variety of downstream tasks and effectively control the output behaviors of LPLMs. Recent studies have proposed numerous methods for fine-tuning a small number of parameters based on open-source LPLMs, reducing the demand for computational and storage resources. Among these, reparameterization fine-tuning methods represented by LoRA (Low-Rank Adaptation) have gained popularity. We find that although these methods perform well in many aspects, there is still considerable room for improvement in terms of complex task adaptability, performance, stability, and algorithm complexity. In response to this, inspired by the idea that the functions of the brain are shaped by its geometric structure, this paper integrates this idea into LoRA technology and proposes a new matrix transformation-based reparameterization method for efficient fine-tuning, named Matrix-Transformation based Low-Rank Adaptation (MTLoRA). The spatiotemporal patterns of brain neural activity are the excitation of different wavelength characteristic patterns of its geometric structure. MTLoRA aims to dynamically alter its spatial geometric structure by applying a transformation-matrix $T$ to perform linear transformations, such as rotation, scaling, and translation, on the task-specific parameter matrix, generating new matrix feature patterns (eigenvectors) to mimic the fundamental influence of complex geometric structure feature patterns in the brain on functions, thereby enhancing the model’s performance in downstream tasks. The transformation-matrix $T$ contains four different structures, each designed to simulate the geometric feature patterns of the brain at different levels. In Natural Language Understanding (NLU) tasks, it is evaluated using the GLUE benchmark test, and the results reveal that MTLoRA achieves an overall performance increase of about 1.0% across eight tasks and reduces the standard deviation by 0.7% in the Corpus of Linguistic Acceptability (CoLA) task; in Natural Language Generation (NLG) tasks, MTLoRA improves performance by an average of 0.95% and 0.56% in the DART and WebNLG tasks, respectively.
@@ 1 Introduction
In this paper, through empirical experiments on two categories of 11 tasks in NLU and NLG, the effectiveness of the MTLoRA method is verified. Compared to the LoRA method, it is evident from the experimental results table 1 that MTLoRA improves performance by about 1.54% ($\sigma$=0.1%) on the CoLA (Warstadt et al., 2019) task, effectively reducing the standard deviation; it enhances performance by about 3.61% ($\sigma$=0.8%) on the RTE (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) task; it increases performance by about 2.45% ($\sigma$=0.3%) on the MRPC (Dolan and Brockett, 2005) task; and it boosts performance by about 0.88% ($\sigma$=0.0%) on the QQP (Quora Question Pairs) task. From the experimental results table 2, it is known that MTLoRA achieves an average performance increase of about 0.95% ($\sigma$=0.1%) on the DART (Nan et al., 2020) task and about 0.56% ($\sigma$=0.1%) on the WebNLG (Gardent et al., 2017) task.
@@ 4.2.3 Experimental Results
$\mathbf{87.49}_{\pm.1}$ $\mathbf{94.5}_{\pm.1}$ $\mathbf{63.11}_{\pm.1}$ $\mathbf{91.25}_{\pm.0}$ $\mathbf{92.6}_{\pm.2}$ $\mathbf{87.73}_{\pm.8}$ $\mathbf{90.69}_{\pm.3}$ $\mathbf{91.37}_{\pm.0}$ $\mathbf{87.34}$
@@
In contrast, the ICFM structure shows robust performance across numerous tasks, particularly excelling in tasks related to semantic similarity and entailment when there is a lack of comprehensive corpus data. For example, it achieves an average performance boost of approximately 1.22% ($\sigma$=0.6%) on the MRPC task and 1.08% ($\sigma$=0.6%) on the RTE task.
@@ 4.3.3 Experimental results
MTLoRA(best) $70.23$ $48.31$ $55.98$
MTLoRA $70.02_{\pm.2}$ $\mathbf{48.05}_{\pm.1}$ $\mathbf{55.86}_{\pm.1}$
@@
MTLoRA-ICFM(best) $70.23$ $47.9$ $55.98$
MTLoRA-ICFM $70.02_{\pm.2}$ $47.75_{\pm.1}$ $55.52_{\pm.5}$
@@
The incorporation of the SHIM structure led to an enhancement in the model’s efficacy on the DART and WebNLG tasks, with an improvement of approximately 0.83% ($\sigma$=0.0%) and 0.56% ($\sigma$=0.1%), respectively. Similarly, the implementation of the ICFM structure resulted in a performance increment of roughly 0.65% ($\sigma$=0.1%) for DART and 0.22% ($\sigma$=0.5%) for WebNLG. The application of the CTCM structure contributed to a performance uplift on the DART and WebNLG tasks by approximately 0.88% ($\sigma$=0.0%) and 0.27% ($\sigma$=0.2%), respectively. Lastly, the adoption of the DTSM structure facilitated a performance boost on the DART and WebNLG tasks by an estimated 0.95% ($\sigma$=0.1%) and 0.31% ($\sigma$=0.3%), respectively.
@@ 4.4 Experimental Analysis
Performance of LoRA on E2E Tasks. To verify whether the MTLoRA method actually reduces the model’s performance on the E2E task in NLG, we re-executed the performance test of LoRA on the E2E task using the same computational environment as MTLoRA. The experimental results show that LoRA achieves an average BLEU score of 69.52% on the E2E task, with a standard deviation of 0.5%. Its average BLEU score is comparable to that of MTLoRA, but with an increased standard deviation. This indicates that the experimental computational environment also has an impact on the results, suggesting that MTLoRA does not actually reduce the model’s performance on the E2E task in NLG.
Sensitivity Analysis of Rank ($r$). This study employs the MTLoRA and LoRA methods based on RoBERTa-base, conducting experiments on the COLA task with varying sizes of rank ($r$) to assess the specific impacts of rank ($r$) on model efficacy and stability. All experiments were carried out with the same random seed, and the specific settings of the experiments are consistent with those described in Section 4.2.2, with the corresponding results displayed in Figure 5:
Figure 5: The MTLoRA and LoRA methods, based on the RoBERTa-base model, were experimentally evaluated on the COLA task using different rank ($r$) sizes. A higher Matthew’s correlation value indicates better model performance.
Several observations can be made from the experimental results in Figure 5: Firstly, the MTLoRA method continues to exhibit good performance and stability on the COLA task across different rank sizes without evident overfitting. Particularly, the ICFM structure in MTLoRA demonstrates exceptional performance and stability, indicating that the ICFM structure can effectively capture the task’s key features, thereby efficiently preventing overfitting. Secondly, compared to the traditional LoRA, MTLoRA can achieve performance improvements with larger ranks while adding fewer parameters. Lastly, the effectiveness of MTLoRA is not solely due to the increased parameters at the same rank, as the performance curve within the examined range suggests that MTLoRA can outperform some higher-ranked LoRAs even at lower ranks.